# 3D VR Sketch Guided 3D Shape Prototyping and Exploration

### v1

@@ 3.3 Conditional shape generation
As shown in Fig. 2, a sparse sketch can represent multiple 3D shapes, generally following the sparse sketch strokes. Therefore, we would like to be able to generate multiple 3D shapes given a sketch input. We achieve this by training a conditional normalizing flow (CNF) model in the latent space.
Specifically, we model the conditional distribution of shape embeddings using a RealNVP network [15] with five layers as in [44]. It transforms the probability distribution of shape feature embedding $p_{e}(\boldsymbol{e}_{f})$ to a unit Gaussian distribution $p_{z}(\boldsymbol{z})$. We obtain the sketch embedding vector $e_{g}^{\prime}$ as described in the previous section, which serves as a condition for our normalizing flow model. Please note that the sketch encoder parameters $\phi$ are frozen at this stage. The sketch condition $(\boldsymbol{e}^{\prime}_{g})^{i}$ is concatenated with the matching 3D shape feature vector $e^{i}_{f}$ at each scale and translation coupling layers, following RealNVP[15]. For convenience, we omit in the following equations a superscript $i$:
$\quad{z}_{i}^{1:d}={e}_{f}^{1:d}\quad\text{and}$ (13)
${z}_{i}^{d+1:D}={e}_{f}^{d+1:D}\odot\exp\left(s\left(\left[e^{\prime}_{g};{e}_{f}^{1:d}\right]\right)\right)+t\left(\left[e^{\prime}_{g};{e}_{f}^{1:d}\right]\right)$ (14)
@@
The idea of the normalizing flow model is to approximate a complicated probability distribution with a simple distribution through a sequence of invertible nonlinear transforms. We train the flow model by maximizing the log-likelihood $\log\left(p_{e}\left({e}\right)\right)$:
$\begin{split}\mathcal{L}_{\text{CNF}}&=-\log\left(p_{e}\left({e}\right)\right)\\ &=-(\log\left(p_{z}\left({z}\right)\right)+\log\left(\left|\operatorname{det}\left(\frac{\partial F\left({e}\right)}{\partial{z}^{T}}\right)\right|\right)),\end{split}$ (15)
where $F(\cdot)$ is the normalizing flow model, and ${\partial F\left({e}\right)}/{\partial{z}^{T}}$ is the Jacobian of $F$ at $e$.
@@
First, the flow module $F$ is updated with the gradients from $\mathcal{L}_{\text{CNF}}$. Then, for each sketch condition $e_{g}^{\prime}$, we randomly sample $K$ different noise vectors $\{z_{k}\}$ from the unit Gaussian distribution $z_{k}\in\mathcal{N}(0,1)$, as shown in Fig. 3 (b). These noise vectors are mapped back to the shape embedding space through the reverse path of the flow model. During training, the obtained shape embeddings $\{e^{\prime\prime}_{g,k}\}$ are fed to the implicit decoder $D_{\theta}(\cdot)$ together with sketch stroke points $g$. Formally, this loss takes the following form:
$\mathcal{L}_{\text{sketch,CNF}}=\frac{\lambda}{N_{s}K}\sum_{i=1}^{N_{s}}\sum_{k=1}^{K}|D(e^{\prime\prime}_{g,k},g_{i})|,$ (16)
@@
During inference, given an input sketch, represented as a set of points $g$, we first obtain its embedding $\boldsymbol{e}_{g}^{\prime}$ using the encoder $E_{\phi}(\cdot)$. We then condition the normalizing flow network with $\boldsymbol{e}_{g}^{\prime}$ and a random noise vector sampled from the unit Gaussian distribution to obtain a shape embedding ${e^{\prime\prime}_{g}}$. We obtain the mean embedding by using the mean of the normal distribution. Finally, this shape embedding is fed into implicit decoder $D_{\theta}(\cdot)$ to obtain a new set of SDF values $\{s^{\prime\prime}|s^{\prime\prime}=D(\tilde{e^{\prime\prime}_{g}})\}$. A 3D geometry is then reconstructed by applying the marching cubes algorithm [31].
@@ 4.1 Implementation Details
During training, for each shape, we sample locations of the 3D points at which we calculate SDF values, as suggested in [40]. We sample two sets of points: close to the shape surface and uniformly sampled in the unit box. Then, the loss is evaluated on the random subsets of those pre-computed points. During inference, the 3D points are sampled on a regular ($256\times 256\times 256$) grid.
@@ 4.2 Datasets
The sketches in the used dataset have a consistent orientation with reference 3D shapes, but might be not well aligned horizontally and vertically to the references, and can have a different scale. We sample shape point clouds and compute SDF values for the normalized 3D shapes as provided in ShapeNetCore-v2, which ensures consistency between the two 3D shape representations. We then normalize the sketches to fit a unit bounding box, following the normalization in the ShapeNetCore-v2. To further improve alignment between sketches and 3D models, we translate sketches, so that their centroids match the centroids of reference shapes.
@@ 4.3 Evaluation Metrics
First, to evaluate the ability of our auto-encoder to faithfully regress 3D shape SDF values given 3D shape point cloud, we can compare the reconstructed 3D shape, denoted as $y_{f}$, with the ground-truth 3D shape, denoted as $y$. In this case, we define shape fidelity measure as $\mathcal{F}_{\text{shape}}(y_{f})=CD(y_{f},y)$.
Then, while the sketches in the used dataset do not align perfectly with reference 3D shapes and contain ambiguity, it is meaningful to expect that the reconstructed 3D shape still should be close to the reference 3D shape. Therefore, we evaluate how close the reconstructed 3D shapes are to the ground-truth when (1) the shape is reconstructed from the sketch embedding $\boldsymbol{e}^{\prime}_{g}$, denoted as $y^{\prime}_{g}$; (2) the shape is reconstructed from the mean conditional embedding of the CNF model, denoted as $\tilde{y}_{g}$; and (3) the shape is reconstructed from a random sample from the latent space of the CNF model, denoted as ${y}^{\prime\prime}_{g}$. The respective losses are: $\mathcal{F}_{\text{shape}}(y^{\prime}_{g})$, $\mathcal{F}_{\text{shape}}(\tilde{y}_{g})$ and $\mathcal{F}^{\text{avg}}_{\text{shape}}({y^{\prime\prime}}_{g,i})=\sum_{i=1}^{5}\mathcal{F}_{\text{shape}}({y}^{\prime\prime}_{g,i})$, where in the latter case we generate $5$ samples and report an average loss value.
@@
Since the used sketches are ambiguous and are not perfectly aligned to a reference, we evaluate the fidelity of the reconstructions to the sketch input, using the loss similar to Eq. 9: $\mathcal{F}_{\text{sketch}}(y)=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}s_{y}(g_{i})$, where $g_{i}$ is the $i$-th sample point from the input sketch and $s_{y}$ denotes the SDF for the shape $y$. Therefore, we also define $\mathcal{F}^{avg}_{\text{sketch}}(y^{\prime\prime}_{g,i})=\frac{1}{5}\sum_{i=1}^{5}\mathcal{F}_{\text{sketch}}(s_{y^{\prime\prime}}(g_{i}))$, as the average fidelity of multiple samples from the CNF model space to an input sketch.
@@ 4.4.1 Deterministic sketch to shape generation
Method Loss $\mathcal{F}_{shape}(y^{\prime}_{g})\downarrow$ $\mathcal{F}_{shape}(y_{f})\downarrow$
@@
Table 1: Evaluation of auto-encoder training strategies with respect to the fidelity ($\mathcal{F}_{shape}(\cdot)$) of the reconstructed 3D shapes to the reference/ground-truth 3D shapes, depending on the used data. Here, $y^{\prime}_{g}$ and $y_{f}$ are reconstructions from an input sketch and 3D shape, respectively. Shape AE, Sketch AE, and Joint AE stand for training encoder only with shape inputs, sketch inputs, or both, respectively.
Our first goal is to learn to map sketches to 3D shapes in a deterministic fashion. One of the challenges in our work comes from the limited dataset size, which is a common factor that should be taken into consideration when working with freehand sketches. Therefore, we proposed training the sketch-to-shape auto-encoder in multiple steps, and in addition, we propose to use a joint auto-encoder, and we use 3D shapes without paired sketches in the NCE loss to improve robustness. Tab. 1 shows that our strategy indeed outperforms alternative strategies. It allows reconstructing 3D shapes similar to reference 3D shapes, as shown by $\mathcal{F}_{shape}(y^{\prime}_{g})$. The fact that $\mathcal{F}_{shape}(y_{f})$ stays low in our proposed design implies that if the sketch is very detailed and accurate, we will obtain careful 3D shape reconstructions.
Method $\mathcal{F}_{shape}(y^{\prime}_{g})\downarrow$ $\mathcal{F}_{shape}(y_{f})\downarrow$
@@ 4.4.2 Conditional sketch to shape generation
Method $K$ $\mathcal{F}_{shape}(y^{\prime}_{g})\downarrow$ $\mathcal{F}^{avg}_{\text{sketch}}(y^{\prime\prime}_{g,i})\downarrow$ $\mathcal{F}^{\text{avg}}_{\text{shape}}(y^{\prime\prime}_{g,i})\downarrow$ $\mathcal{D}_{\text{gnrtns}}\uparrow$
@@
CNF($e_{f}^{\prime}$) - 0.373 0.026$\pm$0.036 0.380 0.043
CNF($e_{f}$) - 0.385 0.030$\pm$0.039 0.420 0.165
CNF($e_{f}$) + $\mathcal{L}_{\text{sketch,CNF}}$ 1 0.366 0.019$\pm$0.034 0.422 0.158
@@
Shape encoding choices.
Note that when training CNF model, we use shape latent code, $\boldsymbol{e}_{f}$, obtained via an inversion process with Eq. 3. Tab. 3, lines 2 and 3, shows that this allows to greatly increase the diversity of the generated results compared to using latent shape codes, $\boldsymbol{e}^{\prime}_{f}$, obtained from the encoder. This comes with a small decrease in fidelity to the reference shape, but this is an expected behavior. The fidelity to the sketch increases just a little bit taking into account standard deviation values of $\mathcal{F}_{\text{sketch}}(e^{\prime\prime}_{g,i})$ in Eq. 3. This result reinforces our design choice of training the auto-decoder first, providing a richer latent space.
@@ 5 Conclusion and Discussion
We present the first method for multiple 3D shape generation conditioned on sparse and abstract sketches. We achieve good fidelity to the input sketch combined with the shape diversity of the generated results. In our work, we show how to efficiently overcome the limitation of the small dataset. Our experiments are currently limited to a single category, but none of the components of our method explicitly exploits any priors about this category. In the future, we would like to extend this work by (1) further improving the input sketch fidelity, potentially taking perceptual multi-view losses into account during training, and (2) considering alternative shape representation for our auto-decoder.
@@
Data $\mathcal{F}_{shape}(y^{\prime}_{g})\downarrow$ $\mathcal{F}^{avg}_{\text{sketch}}(y^{\prime\prime}_{g,i})\downarrow$ $\mathcal{F}^{\text{avg}}_{\text{shape}}(y^{\prime\prime}_{g,i})\downarrow$ $\mathcal{D}_{\text{gnrtns}}\uparrow$
@@
S1.2 Interpolation in sampling space

### v6

@@ 3.3 Conditional shape generation
As shown in Fig. 2, a sparse sketch can represent multiple 3D shapes, generally following the sparse sketch strokes. Therefore, we would like to be able to generate multiple 3D shapes given a sketch. We achieve this by training a conditional normalizing flow (CNF) model in the latent space.
Specifically, we model the conditional distribution of shape embeddings using a RealNVP network [15] with five layers as in [46]. It transforms the probability distribution of shape feature embedding $p_{\boldsymbol{d}}(\boldsymbol{d}^{g})$ to a unit Gaussian distribution $p_{z}(\boldsymbol{z})$. We obtain the sketch embedding vector $\boldsymbol{e}^{c}$ as described in the previous section, which serves as a condition for our normalizing flow model. Please note that the sketch encoder parameters $\phi$ are frozen at this stage. The sketch condition $\boldsymbol{e}^{c}$ is concatenated with the matching 3D shape feature vector $\boldsymbol{d}^{g}$ at each scale and translation coupling layers, following RealNVP [15]:
$\quad{z}^{1:d}={{d}^{g}}^{1:d}\quad\text{and}$ (13)
${z}^{d+1:D}={{d}^{g}}^{d+1:D}\odot\exp\left(s\left(\left[\boldsymbol{e}^{c};{{d}^{g}}^{1:d}\right]\right)\right)+t\left(\left[\boldsymbol{e}^{c};{{d}^{g}}^{1:d}\right]\right)$ (14)
@@
The idea of the normalizing flow model is to approximate a complicated probability distribution with a simple distribution through a sequence of invertible nonlinear transforms. We train the flow model by maximizing the log-likelihood $\log\left(p_{\boldsymbol{d}}\left({\boldsymbol{d}}\right)\right)$:
$\begin{split}\mathcal{L}_{\text{CNF}}&=-\log\left(p_{\boldsymbol{d}}\left(\boldsymbol{d}\right)\right)\\ &=-(\log\left(p_{z}\left({z}\right)\right)+\log\left(\left|\operatorname{det}\left(\frac{\partial F\left(\boldsymbol{d}\right)}{\partial{z}^{T}}\right)\right|\right)),\end{split}$ (15)
where $F(\cdot)$ is the normalizing flow model, and ${\partial F\left(\boldsymbol{d}\right)}/{\partial{z}^{T}}$ is the Jacobian of $F$ at $\boldsymbol{d}$.
@@
First, the flow module $F$ is updated with the gradients from $\mathcal{L}_{\text{CNF}}$. Then, for each sketch condition $\boldsymbol{e}^{c}$, we randomly sample $K$ different noise vectors $\{z_{k}\}$ from the unit Gaussian distribution $z_{k}\in\mathcal{N}(0,1)$, as shown in Fig. 3 (b). These noise vectors are mapped back to the shape embedding space through the reverse path of the flow model. During training, the obtained shape embeddings $\{\boldsymbol{d}^{c}_{k}\}$ are fed to the implicit decoder $D_{\theta}(\cdot)$ together with sketch stroke points $p^{c}_{i}$. Formally, this loss takes the following form:
$\mathcal{L}_{\text{sketch,CNF}}=\frac{\lambda}{N_{s}K}\sum_{i=1}^{N_{s}}\sum_{k=1}^{K}|D\left(\boldsymbol{d}^{c}_{k},p^{c}_{i}\right)|,$ (16)
@@
During inference, given an input sketch, represented as a set of points, we first obtain its embedding $\boldsymbol{e}^{c}$ using the encoder $E_{\phi}(\cdot)$. We then condition the normalizing flow network with $\boldsymbol{e}^{c}$ and a random noise vector sampled from the unit Gaussian distribution to obtain a shape embedding $\boldsymbol{d}^{c}$. We obtain the mean embedding by using the mean of the normal distribution. Finally, this shape embedding is fed into implicit decoder $D_{\theta}(\cdot)$ to obtain a new set of SDF values $\{s^{d^{c}}|s^{d^{c}}=D(\boldsymbol{d}^{c})\}$. A 3D geometry is then reconstructed by applying the marching cubes algorithm [32].
@@ 4.1 Implementation Details
During training, for each shape, we sample locations of the 3D points at which we calculate SDF values. We sample two sets of points: close to the shape surface and uniformly sampled in the unit box. Then, the loss is evaluated on the random subsets of those pre-computed points. During inference, the 3D points are sampled on a regular ($256\times 256\times 256$) grid.
@@ 4.2 Datasets
The sketches in the used dataset have a consistent orientation with reference 3D shapes, but might be not well aligned horizontally and vertically to the references, and can have a different scale. We sample shape point clouds and compute SDF values for the normalized 3D shapes as provided in ShapeNetCore-v2, which ensures consistency between the two 3D shape representations. We then normalize the sketches to fit a unit bounding box, following the normalization in the ShapeNetCore-v2. To further improve alignment between sketches and 3D shapes, we translate sketches, so that their centroids match the centroids of reference shapes.
@@ 4.3 Evaluation Metrics
First, we evaluate the ability of our auto-encoder to faithfully regress 3D shape SDF values given a 3D shape point cloud. We evaluate the fidelity of the regressed 3D shape, $G^{e^{g}}$, to the ground-truth 3D shape, $G$, as follows: $\mathcal{F}_{\text{shape}}(G^{e^{g}})=CD(G,G^{e^{g}})$.
Then, while the sketches in the used dataset do not align perfectly with reference 3D shapes and contain ambiguity, it is meaningful to expect that the reconstructed 3D shape still should be close to the reference 3D shape. Therefore, we evaluate how close the reconstructed 3D shapes are to the ground-truth when (1) the shape is reconstructed from the sketch embedding $\boldsymbol{e}^{c}$, denoted as $G^{{e}^{c}}$; (2) the shape is reconstructed from the predicted conditional mean $\bar{z}^{c}$ of the CNF model, denoted as $G^{\bar{z}^{c}}=D(F^{-1}(\bar{z}^{c}))$; and (3) the shape is reconstructed from a random sample from the latent space of the CNF model, denoted as $G^{d^{c}}$. The respective losses are: $\mathcal{F}_{\text{shape}}(G^{{e}^{c}})$, $\mathcal{F}_{\text{shape}}(G^{\bar{z}^{c}})$ and $\mathcal{F}^{\text{avg}}_{\text{shape}}(G^{d^{c}})=\sum_{i=1}^{5}\mathcal{F}_{\text{shape}}(G^{d^{c}}_{i})$, where in the latter case we generate $5$ samples and report an average loss value.
@@
Since the used sketches are ambiguous and are not perfectly aligned to a reference, we evaluate the fidelity of the reconstructions to the sketch input, using the loss similar to Eq. 9: $\mathcal{F}_{\text{sketch}}(G^{c})=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}s^{c}(p^{c}_{i})$, where $p^{c}_{i}$ is the $i$-th sample point from the input sketch and $s^{c}$ denotes the predicted SDF. With that, we define $\mathcal{F}^{avg}_{\text{sketch}}(s^{d^{c}})=\frac{1}{5}\sum_{i=1}^{5}\mathcal{F}_{\text{sketch}}(s^{d^{c}}_{i})$, as the average fidelity of multiple samples from the CNF model space to an input sketch.
@@ 4.4.1 Deterministic sketch to shape generation
Method Loss $\mathcal{F}_{shape}(G^{{e}^{c}})\downarrow$ $\mathcal{F}_{shape}(G^{e^{g}})\downarrow$
@@
Table 1: Evaluation of auto-encoder training strategies with respect to the fidelity ($\mathcal{F}_{shape}(\cdot)$) of the reconstructed 3D shapes to the reference/ground-truth 3D shapes, depending on the used data. Here, $G^{{e}^{c}}$ and $G^{e^{g}}$ are reconstructions from an input sketch and 3D shape, respectively. Shape AE, Sketch AE, and Joint AE stand for training encoder only with shape inputs, sketch inputs, or both, respectively.
Our first goal is to learn to map sketches to 3D shapes in a deterministic fashion. One of the challenges in our work comes from the limited dataset size, which is a common factor that should be taken into consideration when working with freehand sketches. Therefore, we proposed training the sketch-to-shape auto-encoder in multiple steps, and in addition, we propose to use a joint auto-encoder, and we use 3D shapes without paired sketches in the NCE loss to improve robustness. Tab. 1 shows that our strategy indeed outperforms alternative strategies. It allows reconstructing 3D shapes similar to reference 3D shapes, as shown by $\mathcal{F}_{shape}(G^{{e}^{c}})$. The fact that $\mathcal{F}_{shape}(G^{e^{g}})$ stays low in our proposed design implies that if the sketch is very detailed and accurate, we will obtain careful 3D shape reconstructions.
Method $\mathcal{F}_{shape}(G^{{e}^{c}})\downarrow$ $\mathcal{F}_{shape}(G^{e^{g}})\downarrow$
@@ 4.4.2 Conditional sketch to shape generation
Method $K$ $\mathcal{F}_{shape}(G^{{e}^{c}})\downarrow$ $\mathcal{F}^{avg}_{\text{sketch}}(G^{d^{c}})\downarrow$ $\mathcal{F}^{\text{avg}}_{\text{shape}}(G^{d^{c}})\downarrow$ $\mathcal{D}_{\text{gnrtns}}\uparrow$
@@
CNF($\boldsymbol{e}^{g}$) - 0.373 0.026$\pm$0.036 0.380 0.043
CNF($\boldsymbol{d}^{g}$) - 0.385 0.030$\pm$0.039 0.420 0.165
CNF($\boldsymbol{d}^{g}$) + $\mathcal{L}_{\text{sketch,CNF}}$ 1 0.366 0.019$\pm$0.034 0.422 0.158
@@
Shape encoding choices
Note that when training CNF model, we use the shape latent code, $\boldsymbol{d}^{g}$, obtained via an inversion process with Eq. 3. Tab. 3, lines 2 and 3, shows that this allows to greatly increase the diversity of the generated results compared to using latent shape codes, $\boldsymbol{e}^{g}$, obtained from the encoder. This comes with a small decrease in fidelity to the reference shape, while the fidelity to the sketch increases a little bit. This result reinforces our design choice of training the auto-decoder first, providing a richer latent space.
@@ 5 Conclusion and Discussion
We present the first method for multiple 3D shape generation conditioned on sparse and abstract sketches. We achieve good fidelity to the input sketch combined with the shape diversity of the generated results. In our work, we show how to efficiently overcome the limitation of small datasets. Our experiments are currently limited to a single category, but none of the components of our method explicitly exploits any priors about this category. In the future, we would like to extend this work by (1) further improving the input sketch fidelity, potentially taking perceptual multi-view losses into account during training, and (2) considering alternative shape representation for our auto-decoder.
@@
Data $\mathcal{F}_{shape}(G^{{e}^{c}})\downarrow$ $\mathcal{F}^{avg}_{\text{sketch}}(G^{d^{c}})\downarrow$ $\mathcal{F}^{\text{avg}}_{\text{shape}}(G^{d^{c}})\downarrow$ $\mathcal{D}_{\text{gnrtns}}\uparrow$
@@
S1.2 Comparison with retrieval
We compute our fidelity and diversity metrics on the top 5 retrieval results of [34] with the ground-truth shape in the gallery. Tab. S5 shows that retrieval results have much lower fidelity to sketches and ground-truth shapes. Additionally, note that if there is no shape close to the sketch in the 3D gallery, the retrieval has no chance to succeed.
S1.3 Evaluation on sketches with style not seen during training
Following [35], our test set includes 5 participants those sketches are not used for training. Tab. S5 shows generalization to lower quality 202 VR chair sketches from memory FVRS-M [34].
S1.4 Comparison with 2D sketch-based modeling
We compare with the performance of a sketch-based modeling method, following closely ours, but conditioned on 2D sketches by novices [43]. These sketches match 3D shapes in our training and test sets of VR sketches. We feed all 3 available views to the [20] encoder, which we train with $\mathcal{L}_{\text{align}}$ (Eq.12) and $\mathcal{L}_{\text{SDF}}$ (Eq.5) and leverage our pre-trained decoder, similar to the setting in Tab. 1, lines 3-4, in the main paper. We obtain $\mathcal{F}_{\text{shape}}^{2D}=0.833$ which is much worse than on 3D sketches: $0.437$ (Tab. 1, line 3).
Experiment $\mathcal{F}^{\text{avg}}_{\text{sketch}}(\cdot)\downarrow$ $\mathcal{F}^{\text{avg}}_{\text{shape}}(\cdot)\downarrow$ $\mathcal{D}_{\text{gnrtns}}\uparrow$
Unseen 5 0.019$\pm$0.035 0.460 0.171
FVRS-M 0.030$\pm$0.041 0.431 0.161
Retrieval [34] 0.061$\pm$0.056 1.132 1.138
Table S5: The first two rows show the performance of our method on (1) the set of sketches by 5 participants those sketches were not observed during training – they represent new shapes and new styles; (2) sketches from the FVRS-M dataset [34]. The last row provides numerical evaluation on the quality of the top-5 3D shapes returned by the retrieval method by Luo et al. [34].
S1.5 Interpolation in sampling space